# Microscopic characteristics of SF6 partial discharge induced by a floating linear metal particle

### v1

@@ Abstract
Direct current (DC) gas insulated transmission lines (GILs) possess unique advantages in urban underground power transmission and underwater power transmission applications Magier et al. (2018). In practice, SF6 partial discharge is usually induced due to the presence of metal particles, and existing particle detection and suppression measures have been proven ineffective for DC GILs Li et al. (2022); Zhuang et al. (2023a). Although a considerable amount of experimental results have indicated that the complicated partial discharge induced by metal particles is the main cause of the above issueMa et al. (2018); Sakai et al. (1999a); Donzel et al. (2022); Kim et al. (2016); Merlino and Kim (2006), the microscopic characteristics of related discharge processes and the space charge effect on particles remain unclear.
Gas discharge simulations can resolve spatiotemporal microscopic characteristics of the discharge Beckers et al. (2023); Nae Cho (2013). Chen et al. Chen et al. (2022)simulated metal particle-induced breakdown within a 200 $\upmu$m microgap. Sun et al. Sun et al. (2021)and Zhong et al. Zhong et al. (2022)simulated the impact of the metal particle’s field electron emission on breakdown. However, the discharge gas used in their models is different from that used in GILs. To date, little numerical simulation has been performed on high-pressure pure SF6 Gallimberti and Wiegart (1986); Morrow (1987); Dhali and Pal (1988); Morrow (1991); Espel et al. (2002); Gao et al. (2018); Yang et al. (2022) due to the difficulty of numerical calculations for strongly electronegative gases, and existing results have been reported only for single-gap electrodes, such as electrode protrusions and parallel plates. However, models without relying on the particle-structure scenario cannot obtain the multistage discharge process and interaction between space charge and metal particles, compared to combined gas gap with floating particles in GILs. In particular, under the negative high voltage, linear metal particles often exhibit a firefly motion and significantly reduce the efficiency of diagnosis and suppression Zhuang et al. (2023b); Diessner and Trump (1970). One key determinant of firefly motion is the reversal of the electric field force affected by space charge. Chang et al. Chang et al. (2023) used the point charge model to analyze this force and reported that the polarity of the overall charge $Q$ of a metal particle reverses when firefly motion occurs. However, the surface charge density of linear particles is highly nonuniform and directly determined by the transient interaction between space charge and metal particle, which should be further investigated. To date, a precise understanding of the microscopic characteristics of SF6 discharge induced by floating metal particles is still awaiting.
@@
In this letter, we report the microscopic discharge characteristics (both the discharge propagation and interaction between space charge and metal particle) of SF6 partial discharge induced by a linear metal particle in a coaxial cylinder electrode and reveal the physical mechanism of the metal particle’s electric field force reversal phenomenon, which usually occurs under negative DC high voltage and largely affects the firefly motion. A 2D fluid simulation model is established, including plasma fluid equations Sima et al. (2012) and 18 dominant plasma chemical reactions Zhang et al. (2023); Luo et al. (2020); Ou et al. (2020).
@@
where $\mathbf{n}\cdot\mathbf{D}$ represents the normal components of the electric displacement on the particle surface. The above settings ensure that the electric field on the metal particle surface is normal to the surface and that the entire charge on the metal particle is distributed on the surface. The distribution of surface charge satisfies the electrostatic induction conditions of the metal, which ensures that the metal particle is in a state of electrostatic equilibrium at all time.
A schematic of the discharge system is shown in Fig.1. Here for simplicity, a 2D model is established to describe the circular cross-section of the coaxial cylinder electrode. The 2D description generally underestimates the local field enhancement (the field strength around protrusion), which thus cannot quantitatively capture the 3D characteristics of the partial discharge (e.g., transport of the charged species) but can provide a qualitative prediction of the general discharge mechanisms. For simplicity, we use a reduced model similar to that used in many experimental studies Chang et al. (2023); Wang et al. (2016a); Wu et al. (2019a); Wang et al. (2016b); Sun et al. (2018), it is important to acknowledge that the size of the particle is disproportionately enlarged relative to the entire space; nonetheless, the underlying physical mechanisms captured by the reduced model are, essentially, consistent with those derived from the normal-sized model. The inner radius $r$ is 10 mm and the outer radius $R$ is 30 mm. The radius ratio $r/R$ is close to the optimum value of 1/2.718 Hillhouse and Peterson (1973), which reduces the nonuniformity of the electric field distribution. The linear particle length $L$ is 2 mm, and the particle diameter $a$ is 0.2 mm.
@@
Since the timescale of discharge ($\sim$ ns) is much shorter than that of particle movement ($\sim$ ms), the particle is assumed to be floating in the gap. It locates near the central electrode and the entire gap is divided into two parts, with an upper gap distance d of 1 mm between the particle top tip and central electrode. This distance is close to the experimental condition of firefly motion reported in Ref. Wenger et al. (2019). The gas pressure p is set to atmospheric, and thus this $pd$ range qualitatively ensures the same streamer discharge process as that of higher pressure Pedersen (1970). According to existing experimental results You et al. (2017); Cooke et al. (1977); Endo et al. (1980); Morcos et al. (2000); Wu et al. (2019b), it can be inferred that -25 kV is between the particle lifting voltage and the breakdown voltage. As a result, the central electrode is set as a high voltage direct current (HVDC) electrode at a constant -25 kV, and the initial charge of particle $Q_{\text{0}}$ is set to -113.65 pC, which is calculated using formula Tian et al. (2019); Sakai et al. (2003, 1999b):
@@
The electric field is calculated self-consistently based on the Poisson’s equation, and the charge transport is described by the fluid equation. The discharge is initiated wherever the electric field reaches a prebreakdown threshold. Here, the partial discharge starts from both ends of the particle, where the threshold is satisfied as shown in Fig.2. The lower corona characteristics are similar to those reported by Gao et al. Gao et al. (2018), but the upper streamer characteristics are more complicated. Therefore, this study primarily focuses on analyzing the characteristics of upper streamer discharge.
In contrast to the single SF6 streamer induced by a fixed electrode Seeger and Clemen (2013); Seeger et al. (2008), the upper streamer induced by the floating particle here includes three stages, which might cause aliasing of the diagnosis signal. Stage I represents the primary streamer stage, corresponding to Figs.2(a)–2(c), Stage II represents the secondary streamer stage including an upward and a downward secondary streamer, corresponding to Figs.2(d)–2(e), and Stage III represents the streamer extinction stage, corresponding to Fig.2(f).
During Stage I, since the initial floating potential of the metal particle is -35 kV, the upper streamer is similar to the negative streamer, thus there are two electron peaks positioned at the head and tail of the streamer. Another notable characteristic in the streamer channel is the presence of an electron-deficient region, which is more obvious than that in air Wang et al. (2022a) and SF6 gas mixtures Luo et al. (2020); Ou et al. (2020); Levko and Raja (2023). This is because the positive ion region and the head electron peak together shield the channel field, making it lower than $(E/N)_{\text{cr}}$, subsequently, electrons undergo attachment and are converted into negative ions due to the strong electronegativity of SF6, as shown in Fig.3(d)–3(f).
@@
The negative ion region mentioned above plays a significant role in the recovery of the primary streamer channel field in SF6 as shown in Fig.4(h)–4(i), which is different from that in air Luque et al. (2008); Nijdam et al. (2020). Although this field recovery phenomenon in SF6 has been reported in previous studies Dhali and Pal (1988); Morrow (1991), the underlying mechanism remains unclear. Here, we report a synergistic effect of the negative ion region and head electron peak, which together determine the recovery of the channel field. On one hand, the negative ion region enhances the electric field in the streamer channel, on the other hand, as the head electron peak propagates forward, the shielding effect on the channel field away from it weakens. These two factors work together and lead to the field recovery phenomenon.
The field recovery in Stage I is the immediate cause of the upward secondary streamer in Stage II. As shown in Fig.4(j), when the head electron peak of the primary streamer reaches the HVDC electrode, the field recovers to a level higher than $(E/N)_{\text{cr}}$. As the upward secondary ionizing wave propagates forward, the field inside the secondary channel is once again shielded below $(E/N)_{\text{cr}}$ as shown in Fig.4(k). However, the channel field of the secondary streamer in the air has been reported to remain above its critical value Komuro et al. (2018). This is because the mechanisms of the secondary streamer between SF6 and air are fundamentally different. Specifically, the secondary air streamer noted in the literature occurred either after breakdown or at the rising edge of the pulse voltage, which could introduce a decreasing gas density N caused by intense heating after primary penetration Sigmond (1984), or operate under an increasing applied electric field Wang et al. (2007); Briels et al. (2008). However, the above two factors have little effect here, since the particle-induced partial discharge would not cause the connection of the two poles of the power supply, and the change of gas density N can be ignored. Additionally, the effect of attachment instability which plays a key role in secondary air streamer is relatively small in SF6 Sigmond (1984). Therefore, here the secondary streamer is essentially induced by negative ion accumulation due to the strong electronegativity of SF6.
@@
The symbols of $\sigma_{\text{s}}$·$E_{y}$ at the upper and lower ends of the particle are opposite, resulting in competition between the two ends in determining the vertical force acting on the metal particle as shown in Fig.5(e). In general, the space charge is crucial for the force on the particle. This finding is an important supplement to the existing experimental results Chang et al. (2023) and may offer a more precise understanding of the underlying mechanisms of the previously reported firefly phenomenon in GIL systems Wang et al. (2021).
In summary, the results from this work provide an indepth understanding of the SF6 partial discharge with the presence of the metal particle, which may further offer a reference for addressing issues related to partial discharge signal aliasing and severe particle motion in DC GILs. The microscopic characteristics of multistage SF6 discharge are notably distinct from those in air, which may lead to the aliasing of partial discharge signals. Due to the strong electronegativity of SF6, a negative ion region is formed around the top tip of the metal particle, which is dominated by a positive feedback mechanism. In addition, an electric field recovery phenomenon dominated by the synergistic effect of the negative ion region and the head electron peak is reported. The subsequent upward secondary streamer is dominated by this field recovery and the downward secondary streamer is dominated by the residual space charge. Additionally, we analyze the particle charging process and the reversal of the vertical electric force, which is the dominant factor of firefly motion. The approach in this letter could have a broad application in the study of particle-induced discharge mechanisms. Also note that although the established model can qualitatively describe the general physics, the kinetic characteristics of the charged species (e.g., electrons and ions) and the nonlocal discharge behaviors may not be precisely described by the present fluid model, especially during the initial stage of the discharge, which require further investigations through fully kinetic simulations (e.g., particle model or hybrid model Wang et al. (2022b); Li et al. (2012)).

### v2

@@ Abstract
Direct current (DC) gas insulated transmission lines (GILs) possess unique advantages in urban underground power transmission and underwater power transmission applications. In practice, SF6 partial discharge is usually induced due to the presence of metal particles, and existing particle detection and suppression measures have been proven ineffective for DC GILs Li et al. (2022); Zhuang et al. (2023a). Although a considerable amount of experimental results have indicated that the complicated partial discharge induced by metal particles is the main cause of the above issueMa et al. (2018); Sakai et al. (1999), the microscopic characteristics of related discharge processes and the space charge effect on particles remain unclear.
Gas discharge simulations can resolve spatiotemporal microscopic characteristics of the discharge Beckers et al. (2023). Chen et al. Chen et al. (2022)simulated metal particle-induced breakdown within a 200 $\upmu$m microgap. Sun et al. Sun et al. (2021)and Zhong et al. Zhong et al. (2022)simulated the impact of the metal particle’s field electron emission on breakdown. However, the discharge gas used in their models is different from that used in GILs. To date, simulations of SF6 have been reported primarily for single-gap electrode configurationsGallimberti and Wiegart (1986); Morrow (1987); Dhali and Pal (1988); Morrow (1991); Espel et al. (2002); Gao et al. (2018); Yang et al. (2022), and relating fluid models have typically utilized the local field approximation, which may introduce inaccuracies in capturing non-local characteristics, particularly, the dynamics of charged species near the computational boundary. Moreover, models without relying on the particle-structure scenario cannot obtain the multistage discharge process and interaction between space charge and metal particles, compared to combined gas gap with floating particles in GILs. In particular, under the negative high voltage, linear metal particles often exhibit a firefly motion and significantly reduce the efficiency of diagnosis and suppression Zhuang et al. (2023b); Diessner and Trump (1970). One key determinant of firefly motion is the reversal of the electric field force affected by space charge. Chang et al. Chang et al. (2023) used the point charge model to analyze this force and reported that the polarity of the overall charge $Q$ of a metal particle reverses when firefly motion occurs. However, the surface charge density of linear particles is highly nonuniform and directly determined by the transient interaction between space charge and metal particle, which should be further investigated. To date, a precise understanding of the microscopic characteristics of SF6 discharge induced by floating metal particles is still awaiting.
@@
In this letter, we report the microscopic discharge characteristics (both the discharge propagation and interaction between space charge and metal particle) of SF6 partial discharge induced by a linear metal particle in a coaxial cylinder electrode and reveal the physical mechanism of the metal particle’s electric field force reversal phenomenon, which usually occurs under negative DC high voltage and largely affects the firefly motion. A 2D fluid simulation model using local energy approximation is established, including plasma fluid equations Sima et al. (2012) and 18 dominant plasma chemical reactions Zhang et al. (2023); Luo et al. (2020); Ou et al. (2020).
@@
where $\mathbf{n}\cdot\mathbf{D}$ represents the normal components of the electric displacement on the particle surface. The above settings ensure that the electric field on the metal particle surface is normal to the surface and that the entire charge on the metal particle is distributed on the surface. The distribution of surface charge satisfies the electrostatic induction conditions of the metal, which ensures that the metal particle is in a state of electrostatic equilibrium at all time. The proposed model could reveal detailed mechanisms for the discharge induced by the floating metal particle. By employing the local energy approximation, the accuracy of the microscopic characteristics, particularly the dynamics of charged species near particle tips, can be ensured. The approach and findings presented in this letter could contribute to existing research.
A schematic of the discharge system is shown in Fig.1. Here for simplicity, a 2D model is established to describe the circular cross-section of the coaxial cylinder electrode. The 2D description generally underestimates the local field enhancement (the field strength around protrusion), which thus cannot quantitatively capture the 3D characteristics of the partial discharge (e.g., transport of the charged species) but can provide a qualitative prediction of the general discharge mechanisms. For simplicity, we use a reduced model similar to that used in many experimental studies Chang et al. (2023); Wang et al. (2016); Wu et al. (2019a); Sun et al. (2018), the underlying physical mechanisms captured by the reduced model are consistent with those derived from the normal-sized model. The inner radius $r$ is 10 mm and the outer radius $R$ is 30 mm. The radius ratio $r/R$ is close to the optimum value of 1/2.718 Hillhouse and Peterson (1973), which reduces the nonuniformity of the electric field distribution. The linear particle length $L$ is 2 mm, and the particle diameter $a$ is 0.2 mm.
@@
Since the timescale of discharge ($\sim$ ns) is much shorter than that of particle movement ($\sim$ ms), the particle is assumed to be floating in the gap. It locates near the central electrode and the entire gap is divided into two parts, with an upper gap distance d of 1 mm between the particle top tip and central electrode. This distance is close to the experimental condition of firefly motion reported in Ref. Wenger et al. (2019). The gas pressure p is set to atmospheric, and thus this $pd$ range qualitatively ensures the same streamer discharge process as that of higher pressure Pedersen (1970). According to existing experimental results You et al. (2017); Cooke et al. (1977); Endo et al. (1980); Wu et al. (2019b), it can be inferred that -25 kV is between the particle lifting voltage and the breakdown voltage. As a result, the central electrode is set as a high voltage direct current (HVDC) electrode at a constant -25 kV, and the initial charge of particle $Q_{\text{0}}$ is set to -113.65 pC, which is calculated using formula Tian et al. (2019):
@@
The electric field is calculated self-consistently based on the Poisson’s equation, and the charge transport is described by the fluid equation. The initial electrons are provided by pre-ionization close to reality. The discharge is initiated wherever the electric field reaches a prebreakdown threshold. Here, the partial discharge starts from both ends of the particle, where the threshold is satisfied as shown in Fig.2. The lower corona characteristics are similar to those reported by Gao et al. Gao et al. (2018), but the upper streamer characteristics are more complicated. Therefore, this study primarily focuses on analyzing the characteristics of upper streamer discharge.
In contrast to the single SF6 streamer induced by a fixed electrode Seeger and Clemen (2013), the upper streamer induced by the floating particle here includes three stages, which might cause aliasing of the diagnosis signal. Stage I represents the primary streamer stage, corresponding to Figs.2(a)–2(c), Stage II represents the secondary streamer stage including an upward and a downward secondary streamer, corresponding to Figs.2(d)–2(e), and Stage III represents the streamer extinction stage, corresponding to Fig.2(f).
During Stage I, since the initial floating potential of the metal particle is -35 kV, the upper streamer is similar to the negative streamer, thus there are two electron peaks positioned at the head and tail of the streamer. Another notable characteristic in the streamer channel is the presence of an electron-deficient region, which is more obvious than that in air Wang et al. (2022) and SF6 gas mixtures Luo et al. (2020); Ou et al. (2020); Levko and Raja (2023). This is because the positive ion region and the head electron peak together shield the channel field, making it lower than $(E/N)_{\text{cr}}$, subsequently, electrons undergo attachment and are converted into negative ions due to the strong electronegativity of SF6, as shown in Fig.3(d)–3(f).
@@
The negative ion region mentioned above plays a significant role in the recovery of the primary streamer channel field in SF6 as shown in Fig.4(h)–4(i), which is different from that in air Nijdam et al. (2020). Although this field recovery phenomenon in SF6 has been reported in previous studies Dhali and Pal (1988); Morrow (1991), the underlying mechanism remains unclear. Here, we report a synergistic effect of the negative ion region and head electron peak, which together determine the recovery of the channel field. On one hand, the negative ion region enhances the electric field in the streamer channel, on the other hand, as the head electron peak propagates forward, the shielding effect on the channel field away from it weakens. These two factors work together and lead to the field recovery phenomenon.
The field recovery in Stage I is the immediate cause of the upward secondary streamer in Stage II. As shown in Fig.4(j), when the head electron peak of the primary streamer reaches the HVDC electrode, the field recovers to a level higher than $(E/N)_{\text{cr}}$. As the upward secondary ionizing wave propagates forward, the field inside the secondary channel is once again shielded below $(E/N)_{\text{cr}}$ as shown in Fig.4(k). However, the channel field of the secondary streamer in the air has been reported to remain above its critical value Komuro et al. (2018). This is because the mechanisms of the secondary streamer between SF6 and air are fundamentally different. Specifically, the secondary air streamer noted in the literature occurred either after breakdown or at the rising edge of the pulse voltage, which could introduce a decreasing gas density N caused by intense heating after primary penetration Sigmond (1984), or operate under an increasing applied electric field Briels et al. (2008). However, the above two factors have little effect here, since the particle-induced partial discharge would not cause the connection of the two poles of the power supply, and the change of gas density N can be ignored. Additionally, the effect of attachment instability which plays a key role in secondary air streamer is relatively small in SF6 Sigmond (1984). Therefore, here the secondary streamer is essentially induced by negative ion accumulation due to the strong electronegativity of SF6.
@@
The symbols of $\sigma_{\text{s}}$·$E_{y}$ at the upper and lower ends of the particle are opposite, resulting in competition between the two ends in determining the vertical force acting on the metal particle as shown in Fig.5(e). In general, the space charge is crucial for the force on the particle. This finding is an important supplement to the existing experimental results Chang et al. (2023) and may offer a more precise understanding of the underlying mechanisms of the previously reported firefly phenomenon in GIL systems.
In summary, the results from this work provide an indepth understanding of the SF6 partial discharge with the presence of the metal particle. The microscopic characteristics of multistage SF6 discharge are notably distinct from those in air, which may lead to the aliasing of partial discharge signals. Due to the strong electronegativity of SF6, a negative ion region is formed around the top tip of the metal particle, which is dominated by a positive feedback mechanism. In addition, an electric field recovery phenomenon dominated by the synergistic effect of the negative ion region and the head electron peak is reported. The subsequent upward secondary streamer is dominated by this field recovery and the downward secondary streamer is dominated by the residual space charge. Additionally, we analyze the particle charging process and the reversal of the vertical electric force, which is the dominant factor of firefly motion. The approach in this letter could have a broad application in the study of particle-induced discharge mechanisms. And the results in this letter could provide references for both detection signal and motion suppression in the GIL system. As for detection, relating Ultra High Frequency (UHF) signal can be calculated based on the current signal of discharge, with the resolution of nanosecond time, which can provide a reference for solving UHF signal aliasing. As for suppression, due to the critical role of space charge in electric force on metal particles, an "active ion source" could be a potential approach to regulate the particle movement near the electrode, so as to improve the suppression efficiency.